I didn't see this one or any like it when I was looking for help with this situation:
Suppose you travel north for 35 kilometers then travel east 65 kilometers. How far are you from your starting point?

Answer by stanbon(48551)

You need to understand the Pythagorean equation c^2=a^2+b^2
You need to find "c"; you have a=35 and you have b=65.
So, c = sqrt (35^2 + 65^2)
I don't have my calculator here so I'll leave the computation
in your hands.
cheers,
stan H.